OpenStudy (anonymous):

How do I convert complex numbers into trigonomic functions? 5(cos 53.13* + i sin 53.13*) = 5 cis 53.13*

OpenStudy (anonymous):

complex numbers are not trig functions

OpenStudy (anonymous):

if you want to turn your "trig form" into standard form \(a+bi\) then you need only evaluate you will need a calculator for this one

OpenStudy (anonymous):

looks like you have \(3+4i\) here

OpenStudy (anonymous):

how do i imput the cis into my calculator?

OpenStudy (anonymous):

you don't

OpenStudy (anonymous):

you use \(\cos(53.13)\) and \(\sin(53.13)\)

OpenStudy (anonymous):

I don't?

OpenStudy (anonymous):

ohh, ok

OpenStudy (anonymous):

cis is a notational shortcut

OpenStudy (anonymous):

and evaluate everything before the equal?

OpenStudy (anonymous):

the expression you see on the right is just notational shorthand for the expression on the left

OpenStudy (anonymous):

ohh, ok i get it!

OpenStudy (anonymous):

it stands for "cosine theta plus i sine theta" "cis"

OpenStudy (anonymous):

from that link, is the radius or the angle the trigonomic function?

OpenStudy (anonymous):

5 is the radius, 53.13 is the angle but you knew that already since you had \[5\left(\cos(53.13)+i\sin(53.13)\right)\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

what you get out of this is the part that says "result 3.0001 +3.999 i" which probably means this is supposed to be \(3+4i\) in standard form

OpenStudy (anonymous):

the reason it is not exact is that 53.13 is not exact

OpenStudy (anonymous):

is the trigonomic function "-4.8093+1.36772*i" ?

OpenStudy (anonymous):

ok hold on for a second, i sense some serious confusion "-4.8093+1.36772*i" is not a trigonometric function at all. it is a complex number

OpenStudy (anonymous):

\(\sin(x), \cos(x), \tan(x)\) are trig functions \(a+bi\) or \(r\left(\cos(\theta)+i\sin(\theta)\right)\) are numbers

OpenStudy (anonymous):

and so \[5\left(\cos(53.13)+i\sin(53.13)\right)\] is just a number. it is a complex number, but it is a number

OpenStudy (anonymous):

how do i change it to trigonomic form?

OpenStudy (anonymous):

and in fact, it is the number \(3+4i\)

OpenStudy (anonymous):

\[5\left(\cos(53.13)+i\sin(53.13)\right)\] is already in trig form.

OpenStudy (anonymous):

huh so the question is the answer?

OpenStudy (anonymous):

\(r(\cos(\theta)+i\sin(\theta))\) is trig form \(a+bi\) is standard form

OpenStudy (anonymous):

so that means that the equation i posted first is the trigonomic form of that equation?

OpenStudy (anonymous):

i cannot imagine the question was "express \(5\left(\cos(53.13)+i\sin(53.13)\right)\) in trig form" as it is already in trig form but i could be wrong the question might be "express \(3+4i\)in trig form" or "express \(5\left(\cos(53.13)+i\sin(53.13)\right)\) in standard form"

OpenStudy (anonymous):

yes, it is already in trig form.

OpenStudy (anonymous):

the question was to change the complex number in a previous question to trig form, and the answer to the other question was in trig form already

OpenStudy (anonymous):

so i guess.........its the answer for this question too

OpenStudy (anonymous):

what was the exact question?

OpenStudy (anonymous):

"Use your graphing calculator to convert he complex number to trigonomic form in the problems below: then it said the question was #47 #47 is: Write each complex number in trogonomic form. Round all angles to the hundreth os a degree" 3 + 4i

OpenStudy (anonymous):

aaaah so that was the question, wrote \(3+4i\) in trig form, and the answer was \(5\left(\cos(53.13)+i\sin(53.13)\right)\) good enough

OpenStudy (anonymous):

yes, that is the answer for sure

OpenStudy (anonymous):

and then that was also the answer for 55

OpenStudy (anonymous):

lol, two answers for the price of one

OpenStudy (anonymous):

so maybe the question is, how do you write \(3+4i\) in trig form?

OpenStudy (anonymous):

do you know how to do that?

OpenStudy (anonymous):

umm, actually, no

OpenStudy (anonymous):

well that is what you need to turn \(3+4i\) in to \(5\left(\cos(53.13)+i\sin(53.13)\right)\)

OpenStudy (anonymous):

so, in the end, the final answer is 5(cos(53.13)+isin(53.13)) ?

OpenStudy (anonymous):

you need two numbers, \(r\) and \(\theta\) to turn \(a+bi\) in to \(r\left(\cos(\theta)+i\sin(\theta)\right)\)

OpenStudy (anonymous):

yes that is the answer, but i sense that it is not clear how to get there

OpenStudy (anonymous):

r = 5, and theta = 53.13

OpenStudy (anonymous):

yea i know, but do you know where these numbers come from?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

r = square root of : 3^2 + 4^2 = square root of 25

OpenStudy (anonymous):

ok then you are done

OpenStudy (anonymous):

ok, thanks for helping me out! I was totally confused!